# TennisTV: Do Multimodal Large Language Models Understand Tennis Rallies?

### v3

@@ Abstract
Multimodal large language models (MLLMs) excel at general video understanding but struggle with fast, high-frequency sports like tennis, where rally clips are short yet information-dense. To systematically evaluate MLLMs in this challenging domain, we present TennisTV, the first and most comprehensive benchmark for tennis video understanding. TennisTV models each rally as a temporal-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation. It covers 8 tasks from the stroke level to the rally level and includes 2527 human-verified questions. Evaluating 17 representative MLLMs, we provide the first systematic assessment of tennis video understanding. Results yield two key insights: (i) frame-sampling density should be tailored and balanced across tasks, and (ii) improving temporal grounding is essential for stronger reasoning.
@@ 1 Introduction
In order to systematically evaluate the current MLLMs’ understanding of short, information-dense sports videos, as well as promote the community’s improvement of MLLMs’ understanding of tennis videos, we introduce TennisTV, the first high-quality benchmark specifically designed for understanding tennis rally videos. TennisTV defines a stroke event as the atomic unit, spanning movement initiation, stroke action choice, directional execution, and the ball’s landing. Each rally video is modeled as a time-ordered rally sequence composed of these consecutive stroke events. A total of 9 evaluation tasks that match the sports characteristics of tennis, including two types of multiple-choice questions and judgment statements. We design an automated annotation pipeline. Through three-stage annotation process, we realize structured annotation integration for the entire rally based on the rough annotation provided by the original data source[15], extract additional athletes’ movement information, and generate questions. Finally, through manual quality control, we construct the TennisTV benchmark, which contains 2527 curated questions.
We conduct a comprehensive evaluation of 17 state of the art multimodal large language models (MLLMs) on TennisTV, including three closed-source commercial models, eight open-source non-thinking VideoLLMs, and five open-source thinking VideoLLMs. The study yields three findings: first, certain tasks exhibit a high degree of correlation; second, the optimal frame sampling density differs across tasks; third, explicit reasoning delivers gains on knowledge light tasks. From these insights above, we draw two practical takeaways: balance frame sampling density with task granularity, and to further enhance the temporal grounding capabilities of reasoning models. We hope that TennisTV and these insights will guide future advances in tennis video understanding.
@@ 2.1 Sports Understanding
Sports understanding [18] is a rapidly evolving field that encompasses multiple research topics and integrates diverse modalities, covering tasks such as action recognition [3], athlete analysis [17], tactical planning [14], foul recognition [4]. Recent work on MLLMs [9, 22, 21] has shifted toward general frameworks spanning diverse sports tasks, while existing work[15] in tennis primarily uses conventional neural networks for stroke label prediction, with limited modeling of rally level context and causal relations. Consequently, MLLM capabilities in tennis video remain largely unexplored. This paper uses tennis as the empirical domain and proposes a novel benchmark for fast-paced, high-frequency sports video understanding, with the goal of advancing future research in this area.
@@ 4.3 Main Results
Finding I — High degree of correlation among the AR, TI, and TP tasks. Fig 4 reveals a robust positive correlation among AR, TI, and TP that persists across diverse model families and sampling protocols. We attribute this to the fact that these tasks form a coherent capability hierarchy: recognizing specific actions yields the primitives for technique identification, which subsequently supports the inference of tactic preferences. Consequently, performance on one task is highly predictive of the others, suggesting that low-level refinements in attribution or temporal stitching can transfer to tactic inference. Taken together, these findings underscore the empirical soundness of our benchmark.
@@
Finding II — The optimal frame sampling density varies across perceptual levels. In fast-paced, high-frequency sports videos, the number of frames required varies across different levels of understanding. As shown in Table 2, performance with Qwen2.5-VL-7B peaks at 32 frames, while adding more frames yields diminishing or even negative returns, particularly for rally-sequence metrics. These results suggest that finer temporal resolution is most beneficial for capturing brief, high-frequency stroke cues, whereas rally level reasoning saturates earlier and is more sensitive to redundancy.
@@ 5 Takeaways
  • Enhancing the temporal understanding ability of video reasoning models. The experimental results have proved that reasoning can improve rally level understanding, but it has not benefited from fine-grained temporal grounding task. This suggests that future RL post-training should emphasize temporal understanding and alignment mechanisms rather than relying solely on general video reasoning.

### v4

@@ Abstract
Multimodal large language models (MLLMs) excel at general video understanding but struggle with fast, high-frequency sports like tennis, where rally clips are short yet information-dense. To systematically evaluate MLLMs in this challenging domain, we present TennisTV, the first and most comprehensive benchmark for tennis video understanding. TennisTV models each rally as a temporal-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation. It covers 8 tasks from the stroke-level to the rally-level and includes 2527 human-verified questions. Evaluating 17 representative MLLMs, we provide the first systematic assessment of tennis video understanding. Results yield two key insights: (i) frame-sampling density should be tailored and balanced across tasks, and (ii) improving temporal grounding is essential for stronger reasoning.
@@ 1 Introduction
In order to systematically evaluate the current MLLMs’ understanding of short, information-dense sports videos, as well as promote the community’s improvement of MLLMs’ understanding of tennis videos, we introduce TennisTV, the first high-quality benchmark specifically designed for understanding tennis rally videos. TennisTV defines a stroke event as the atomic unit, spanning movement initiation, stroke action choice, directional execution, and the ball’s landing. Each rally video is modeled as a time-ordered rally sequence composed of these consecutive stroke events. A total of 8 evaluation tasks that match the sports characteristics of tennis, including two types of multiple-choice questions and judgment statements. We design an automated annotation pipeline. Through three-stage annotation process, we realize structured annotation integration for the entire rally based on the rough annotation provided by the original data source[15], extract additional athletes’ movement information, and generate questions. Finally, through manual quality control, we construct the TennisTV benchmark, which contains 2527 curated questions.
We conduct a comprehensive evaluation of 17 state of the art MLLMs on TennisTV, including three closed-source commercial models, eight open-source non-thinking VideoLLMs, and five open-source thinking VideoLLMs. The study yields three findings: first, certain tasks exhibit a high degree of correlation; second, the optimal frame sampling density differs across tasks; third, explicit reasoning delivers gains on knowledge light tasks. From these insights above, we draw two practical takeaways: balance frame sampling density with task granularity, and to further enhance the temporal grounding capabilities of reasoning models. We hope that TennisTV and these insights will guide future advances in tennis video understanding.
@@ 2.1 Sports Understanding
Sports understanding [18] is a rapidly evolving field that encompasses multiple research topics and integrates diverse modalities, covering tasks such as action recognition [3], athlete analysis [17], tactical planning [14], foul recognition [4]. Recent work on MLLMs [9, 22, 21] has shifted toward general frameworks spanning diverse sports tasks, while existing work[15] in tennis primarily uses conventional neural networks for stroke label prediction, with limited modeling of rally-level context and causal relations. Consequently, MLLM capabilities in tennis video remain largely unexplored. This paper uses tennis as the empirical domain and proposes a novel benchmark for fast-paced, high-frequency sports video understanding, with the goal of advancing future research in this area.
@@ 4.3 Main Results
Finding I — High degree of correlation among the AR, TI, and TP tasks. Fig 4 reveals a robust positive correlation among AR, TI, and TP that persists across diverse model families and sampling protocols. We attribute this to the fact that these tasks form a coherent capability hierarchy: recognizing specific actions yields the primitives for technique identification, which subsequently supports the inference of tactic preferences. Consequently, performance on one task is highly predictive of the others, suggesting that low-level refinements in attribution or temporal stitching can transfer to tactic inference. Taken together, these outcomes underscore the empirical soundness of our benchmark.
@@
Finding II — The optimal frame sampling density varies across perceptual levels. In fast-paced, high-frequency sports videos, the number of frames required varies across different levels of understanding. As shown in Table 2, performance with Qwen2.5-VL-7B peaks at 32 frames, while adding more frames yields diminishing or even negative returns, particularly for rally-sequence metrics. These results suggest that finer temporal resolution is most beneficial for capturing brief, high-frequency stroke cues, whereas rally-level reasoning saturates earlier and is more sensitive to redundancy.
@@ 5 Takeaways
  • Enhancing the temporal understanding ability of video reasoning models. The experimental results have proved that reasoning can improve rally-level understanding, but it has not benefited from fine-grained temporal grounding task. This suggests that future RL post-training should emphasize temporal understanding and alignment mechanisms rather than relying solely on general video reasoning.